OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x – y = 7 x = y + 3

OpenStudy (anonymous):

A. (y + 3) – y = 7 B. x = (7 – x) + 3 C. x – (y + 3) = 7 D. x – y = y + 3

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

If you plug the second equation into the first you get A xD

OpenStudy (anonymous):

o ty

OpenStudy (anonymous):

A it is than